OpenStudy (anonymous):

Find the average value of f(x) = sin(x) over the interval [0, π].

OpenStudy (anonymous):

https://gyazo.com/d975c3d9750e0048a4741d5d7ead2c04

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

M = average value of f(x) from x = a to x = b \[\Large M = \frac{1}{b-a}\int_{a}^{b}f(x)dx\]

jimthompson5910 (jim_thompson5910):

In this case, a = 0 b = pi f(x) = sin(x) \[\Large M = \frac{1}{b-a}\int_{a}^{b}f(x)dx\] \[\Large M = \frac{1}{\pi-0}\int_{0}^{\pi}\sin(x)dx\] \[\Large M = ??\]

OpenStudy (anonymous):

That's basically what I did without multiplying by 1/pi

OpenStudy (anonymous):

So it should be 2/pi

jimthompson5910 (jim_thompson5910):

you are correct